# Relativistic heavy-ion collisions; classical Yang-Mills theory; glasma

### 09 Jan 2024 v1

@@ 3.3 Broader rapidity region
Figure 4: The transverse and longitudinal pressure in the rapidity region, $-2\leq\eta\leq 2$ for $Q_{\rm s}=1,2,3,4$ and $5$. The upper left and right panels show the transverse and longitudinal pressure for $l_{\rm L}/r_{\rm L}=2$, respectively. The lower left and right panels show the transverse and longitudinal pressure for $l_{\rm L}/r_{\rm L}=0.01$, respectively. The plots with different colors show the results for the different proper times, $Q_{\rm s}\tau=1,2,3,4$ and $5$.
To deepen the understanding of the pressure isotropization in the $3D$ glasma, we examine the pressure results across a broader rapidity region. The figure 4 provides both transverse and longitudinal pressures normalized by the energy density across the broader rapidity region, $-2\leq\eta\leq 2$. The upper left and right panels display the transverse and longitudinal pressures for $l_{\rm L}=2r_{\rm L}$ respectively, and the lower panels showing results for $l=0.01r_{\rm L}$ in a similar arrangement. The plots with different colors show the results for the different proper times, $Q_{\rm s}\tau=1,2,3,4$ and $5$. It is found that around $\eta=0$, both transverse and longitudinal pressures over the energy density are relatively flat, undergoing drastic changes in a larger rapidity range. To understand what contributes to the drastic change in behavior, we consider the evolution of the classical color current by examining its amplitude, defined as
@@
It is found in Fig. 5 that $\rho_{\rm amp.}$ vanishes in the rapidity region for flat $P_{\perp}/\epsilon$ and $P_{\rm L}/\epsilon$, shown in Fig. 4. On the other hand, in the larger rapidity region, where $P_{\perp}/\epsilon$ and $P_{\rm L}/\epsilon$ changes largely, $\rho_{\rm amp.}$ is also found to increase significantly. This correspondence indicates that the former and latter regions should be identified as the glasma and the mixture of the glasma and Weizsäcker-Williams field. Therefore, the flat region in Fig. 4 indicates that the pressure isotropization of the glasma at non-zero rapidity is almost the same as those at mid-rapidity.
Figure 5: The amplitude of the color charge density, defined in Eq. (38), in the rapidity region, $-2\leq\eta\leq 2$ for $Q_{\rm s}=1,2,3,4$ and $5$. The left and right panels show the results for $l_{\rm L}/r_{\rm L}=2$ and $0.01$, respectively. The results are normalized by the saturation scale and coupling constant.
@@ 4 Summary
We have investigated the effects of boost invariance breaking on the isotropization of pressure in the glasma. This breaking is caused by spatial fluctuations of the classical color charge density along the collision axis. Our calculation is done using the $3+1$ D glasma simulation that we have recently developed. In this method, the longitudinal spatial fluctuations is taken into account through the longitudinal correlation length in the color charge density, a model parameter. While the physical origin of this correlation is still a subject of ongoing research, simulations with various correlation lengths offer useful insights into both the pressure evolution of the glasma and a deeper understanding of the $3$D glasma model itself.
We have provided numerical results for pressure and energy density at mid-rapidity and in a broader rapidity region. Our findings show that, despite variations in longitudinal correlation lengths, the evolutions of the pressure isotropy observed are qualitatively similar. At $Q_{\rm s}\tau=0$, the transverse pressure is minimal compared to longitudinal pressure, indicative of the dominance of the Weizsäcker-Williams field. Then, at the early time, the positive transverse pressure and negative longitudinal pressure appear, which indicates the formation of strong longitudinal color electromagnetic fields between separating nuclei like the boost invariant glasma. Finally, the transverse pressure becomes almost half of the energy density, and longitudinal pressure becomes negligible at the late time. Further, it is found that, at this late time, the glasma’s energy density at mid-rapidity decreases as $1/\tau$. Both the negligible longitudinal pressure and the $1/\tau$ decay of energy density in the late stage are consistent with the system expanding longitudinally like a dilute gas, deviating significantly from the Bjorken hydrodynamics with the ideal gas equation of state. By examining the pressure results across a broader rapidity region, we further enhance the understanding of pressure isotropization. The pressure isotropization of the glasma in the non-zero rapidity region is almost the same as that observed in mid-rapidity.
@@
Conceptualization, H.M. and X.-G H.; methodology, H.M.; software, H.M; validation, H.M. and X.-G H.; formal analysis, H.M.; investigation, H.M. and X.-G H.; data curation, H.M.; writing—original draft preparation, H.M. and X.-G H.; writing—review and editing, H.M. and X.-G H.; funding acquisition, X.-G H. All authors have read and agreed to the published version of the manuscript.
@@
Data are available upon request

### 29 Feb 2024 v3

@@ 3.3 Broader Rapidity Region
To deepen the understanding of the pressure isotropization in the $3$D glasma, we examine the pressure results across a broader rapidity region. The Figure 4 provides both transverse and longitudinal pressures normalized by the energy density across the broader rapidity region, $-2\leq\eta\leq 2$. The upper left and right panels display the transverse and longitudinal pressures for $l_{\rm L}=2r_{\rm L}$, respectively, and the lower panels show results for $l=0.01r_{\rm L}$ in a similar arrangement. The plots with different colors show the results for the different proper times, $Q_{\rm s}\tau=1,2,3,4$, and $5$. It is found that around $\eta=0$ both transverse and longitudinal pressures over the energy density are relatively flat, undergoing drastic changes in a larger rapidity range. To understand what contributes to the drastic change in behavior, we consider the evolution of the classical color current by examining its amplitude, defined as
@@
It can be seen in Figure 5 that $\rho_{\rm amp.}$ vanishes in the rapidity region for flat $P_{\perp}/\epsilon$ and $P_{\rm L}/\epsilon$, shown in Figure 4. On the other hand, in the larger rapidity region, where $P_{\perp}/\epsilon$ and $P_{\rm L}/\epsilon$ change significantly, $\rho_{\rm amp.}$ is also found to increase significantly. This correspondence indicates that the former and latter regions should be identified as the glasma and the mixture of the glasma and Weizsäcker–Williams field. Therefore, the flat region in Figure 4 indicates that the pressure isotropization of the glasma at non-zero rapidity is almost the same as those at mid-rapidity.
Figure 4: The transverse and longitudinal pressures in the rapidity region, $-2\leq\eta\leq 2$, for $Q_{\rm s}=1,2,3,4$, and $5$. The upper left and right panels show the transverse and longitudinal pressures for $l_{\rm L}/r_{\rm L}=2$, respectively. The lower left and right panels show the transverse and longitudinal pressures for $l_{\rm L}/r_{\rm L}=0.01$, respectively. The plots with different colors show the results for the different proper times, $Q_{\rm s}\tau=1,2,3,4$, and $5$.
Figure 5: The amplitude of the color charge density, defined in Equation (38), in the rapidity region, $-2\leq\eta\leq 2$, for $Q_{\rm s}=1,2,3,4$, and $5$. The left and right panels show the results for $l_{\rm L}/r_{\rm L}=2$ and $0.01$, respectively. The results are normalized by the saturation scale and coupling constant.
@@ 4 Summary
We have investigated the effects of boost invariance breaking on the isotropization of pressure in the glasma. This breaking is caused by spatial fluctuations of the classical color charge density along the collision axis. Our calculation is performed using the 3+1D glasma simulation that we have recently developed. In this method, the longitudinal spatial fluctuations are taken into account through the longitudinal correlation length in the color charge density, a model parameter. While the physical origin of this correlation is still a subject of ongoing research, simulations with various correlation lengths offer useful insights into both the pressure evolution of the glasma and a deeper understanding of the $3$D glasma model itself.
We have provided numerical results for pressure and energy density at mid-rapidity and in a broader rapidity region. Our findings show that, despite variations in longitudinal correlation lengths, the evolutions of the pressure isotropy observed are qualitatively similar. At $Q_{\rm s}\tau=0$, the transverse pressure is minimal compared to longitudinal pressure, indicative of the dominance of the Weizsäcker–Williams field. Then, at the early time, the positive transverse pressure and negative longitudinal pressure appear, which indicates the formation of strong longitudinal color electromagnetic fields between separating nuclei like the boost invariant glasma. Finally, the transverse pressure becomes almost half of the energy density, and longitudinal pressure becomes negligible at the late time. Further, it is found that, at this late time, the glasma’s energy density at mid-rapidity decreases as $1/\tau$. Both the negligible longitudinal pressure and the $1/\tau$ decay of energy density in the late stage are consistent with the system expanding longitudinally like a dilute gas, deviating significantly from the Bjorken hydrodynamics with the ideal gas equation of state. By examining the pressure results across a broader rapidity region, we further enhance the understanding of pressure isotropization. The pressure isotropization of the glasma in the non-zero rapidity region is almost the same as that observed in mid-rapidity.
@@
We discuss the interpretation of the longitudinal correlation length in our 3D nucleus model. It is roughly estimated as the inverse of the typical longitudinal momentum of the hard partons contributing to the color charge density. This length may be associated with the Bjorken variable $x$, which in turn implies that the correlation length is also related to the saturation scale $Q_{s}(x)$. Addressing the challenge of advancing beyond a basic Gaussian model to more precisely depict the 3D color charge density for a specific Bjorken x is important. The JIMWLK equation offers a framework for calculating observables from the color charge density over a wide range of Bjorken variable $x$ but assumes an extremely thin spatial distribution of color charge density. Then, to accurately determine the spatial rapidity dependence of an observable, a conversion from momentum rapidity to spatial rapidity is essential. Therefore, the feasibility of constructing a spatial 3D color charge density based on the JIMWLK equation presents a complex challenge.
Finally, we point out the gap between the current $3$D glasma model and phenomenology. To enhance the realism of the initial conditions, it would be advantageous to utilize the methodology from the $2$D IP-glasma model IPglasma1 ; IPglasma2 , a phenomenological model of the glasma in the mid-rapidity. This approach allow us to account for the event-by-event fluctuation of nucleon positions, the saturation scale based on the IP-sat model IPsat1 ; IPsat2 , and the expansion of the system in the transverse plane, all of which are not taken into account in the current model. For the transition of glasma into initial hydrodynamic conditions, separating the glasma from the Weizsäcker-Williams(WW) field is crucial. One proposed method is to employ new “usual Milne coordinates” for observation, setting the proper time to 00 at the nucleus’s rear surface, unlike the current usual Milne coordinates, which link proper time of 0 with the nucleus’s central longitudinal position. This alteration would place the majority of the WW field outside the forward light cone, facilitating the separation from the glasma. Considering the energy density and pressure in the local rest frame, defined as eigenvalues of the energy-momentum tensor, could potentially exclude the contribution of the WW field on the EM tensor 3Dglasma7 . Since the EM tensor’s determinant for the WW field is zero, energy density and pressure in the local rest frame should effectively disappear in regions heavily influenced by the WW field. To further align the 3D glasma model with real-world phenomena, transitioning from SU(2) to SU(3) is necessary.
@@
Conceptualization, H.M. and X.-G.H.; methodology, H.M.; software, H.M.; validation, H.M. and X.-G.H.; formal analysis, H.M.; investigation, H.M. and X.-G.H.; data curation, H.M.; writing—original draft preparation, H.M. and X.-G.H.; writing—review and editing, H.M. and X.-G.H.; funding acquisition, X.-G.H. All authors have read and agreed to the published version of the manuscript.
@@
Data are available upon request.
Acknowledgements.